1. Calculate the pH of a mixture of 0.10 M acetic acid and 0.20 M sodium acetate. The pKa of acetic acid is 4.76.

2. Calculate the ratio of acetate and acetic concentrations required to prepare a pH = 5.30 buffer.
Answer #1

[CH3COOH]   = 0.10M

[CH3COONa]   = 0.20M

PKa   = 4.76

PH    = Pka + log[CH3COONa]/[CH3COOH]

        = 4.76 + log0.2/0.1

        = 4.76 + 0.3010

         = 5.061 >>>>>answer

2.

pH = 5.30

Pka = 4.76

PH    = Pka + log[CH3COONa]/[CH3COOH]

5.30 = 4.76 + log[CH3COONa]/[CH3COOH]

log[CH3COONa]/[CH3COOH]    = 5.30-4.76

log[CH3COONa]/[CH3COOH]    = 0.54

[CH3COONa]/[CH3COOH]     = 3.4674

[CH3COONa]:[CH3COOH]      = 3.4674 : 1